# OpticalAging: Real-time Presbyopia Simulation for Inclusive Design via Tunable Lenses

### v1

@@ 3.1. Replication and Generalizability
In developing and evaluating the tunable lens eyewear prototype presented in this paper, we have adhered to a strict confidentiality framework as stipulated by a Non-Disclosure Agreement (NDA) with the device’s providing company, a for-profit enterprise. However, the core parts, the tunable lenses, are commercially available and intensively manufactured (Opt, tune; cor, nses; sin, nics). Regarding academic explorations, Hasan et al. (Hasan et al., 2017) and Li et al. (Li et al., 2006) as well as other related pioneering works (Chen et al., 2023; He et al., 2019; Liu et al., 2023; Agarwala et al., 2022; Milton et al., 2014; Jiang et al., 2021) have detailed the manufacturing process and the core algorithm of adjusting focal-tunable lenses. Our described method is not tied to any specific product and can be generalized to similar devices. Hence, following the underlying principles of our method along with the previously discussed literature, we believe that the optical presbyopia simulation presented in this research can be easily reproduced for future investigations. The comprehensive framework supports the algorithm’s logical workflow and describes the implementation techniques in a manner that is adaptable to various hardware configurations, facilitating replication and modification in diverse research and development environments.
@@ 4. User Study
To explore and validate our approach, our user study was structured into three distinct parts. Part I involved the evaluation of optically shifted near points for simulated presbyopic age groups corresponding to the 40s, 50s, and 60s. Part II required participants to perform 20-minute daily reading activities with our simulation eyewear in the experimental environment. Part III engaged a professional product designer and a design student in a field study to integrate presbyopia-friendly elements into their design workflows, examining the practical applications of our simulation in professional practice. Our research protocols were rigorously reviewed and have obtained approval from the Institutional Review Board (IRB) of our university.
@@
Calibration and Individual Adjustment. The user study began with a careful calibration procedure to tailor to each participant’s unique visual needs. Using tunable lenses’ ability to simulate concave or convex lenses, nearsighted or farsighted subjects participated in the study without their regular corrective eyewear. The corrective values determined by the participants were then incorporated into our simulation algorithm as the offset.
First, the participant puts on the ViXion01 eyewear and adjusts the interpupillary distance to position the tunable lens modules precisely in front of their eyes; see Figure 2 for a detailed hardware structure. Following ViXion01’s official guidelines, and our calibration workflow, see Figure 4, we ensure that the embedded tunable lenses initially functioned as corrective lenses to counteract potential nearsightedness or farsightedness. If the wearer has normal vision, the tunable lenses function as 0D lenses so that they can see things clearly as usual, when the focal plane is far away from the preset near point range. Participants who reported having no noticeable visual acuity difference between two eyes are instructed to manipulate the adjustable dial, which changes the lenses’ diopters until clarity is achieved for a standing hardcover book at a distance of one meter. Individual lens adjustment is also performed for participants reporting anisometropia (disparate visual acuity between their eyes), Figure 4 demonstrates the detailed calibration workflow. In this individual adjustment case, all calibration steps were remotely monitored and controlled using ViXion01’s official smartphone application, the graphical user interface is illustrated in Figure 3 B.
@@ 4.1. Part I: Evaluation of Optically Shifted Near Points
Participants: We recruited 19 individuals from diverse professional backgrounds for this study via word of mouth. They ranged from 18 (with consent from the individual and their guardian) to 35 years old, with an average age (MEAN) of 27.8 years and a standard deviation (SD) of 4.1. We specifically targeted individuals under 40 years old, as those in their 40s typically begin to experience presbyopia (Duane, 1922). All participants were provided with essential information about this user study through a consent form. They took part in the study after signing the consent form. As for the baseline eyesight criteria, because the ViXion01’s built-in tunable lenses can optically correct both nearsightedness and farsightedness, participants were allowed to participate without wearing corrective glasses or contact lenses. Detailed calibration regarding the fitting of their interpupillary distance and corrective focal power is described in the previous Approach section (3). Participants received a 1000 JPY Amazon gift card as compensation for their time.
@@
[…]
Part I: Results
The results of our study on optically shifted near points of accommodation for simulated age groups are presented in Figure 7. This boxplot illustrates the distribution of near point measurements across different presbyopia simulation modes, including the baseline and simulated 40s, 50s, and 60s conditions.
Our findings demonstrate a clear progression in these accommodation near point distances as the simulated age increases: (1) The median near point for participants without simulation was 132 mm, with a range from 61 mm to 231 mm (IQR: 101-162 mm). (2) 40s Simulation: When simulating presbyopia for individuals in their 40s, we observed a substantial shift in the near point. The median increased to 233 mm, with a range from 221 mm to 242 mm (IQR: 228-238 mm). (3) 50s Simulation: For the 50s simulation, we noted a further increase in the near point distance. The median rose to 378 mm, with a range from 356 mm to 396 mm (IQR: 370-382 mm). (4) 60s Simulation: The 60s simulation showed the most dramatic shift in near point distance. The median increased to 782 mm, with a range from 762 mm to 793 mm (IQR: 779-789 mm).
[…]
4.2. Part I: Discussion
The results of our study demonstrate the robust effectiveness of our optical simulation method in shifting the near point to the presbyopic 40s, 50s, and 60s modes. The clear progression in measured near point distances from baseline (median 125 mm) through simulations of 40s (233 mm), 50s (378 mm), and 60s (782 mm) aligns closely with the algorithmically set patterns (Table 2) of presbyopia development (Duane, 1922; Presbyopia, 2016; Schwartz, 2013). This progression was not only visually apparent in our boxplot (Figure 7) but also confirmed by our repeated measures ANOVA ($F(0.78,14.04)=2104.7,p<.001,\eta^{2}=0.992$).
The extremely high effect size ($\eta^{2}=0.992$) underscores the method’s ability to produce a highly controlled, consistent, and substantial simulation of presbyopic near point changes, validating the core mechanism of our approach. This consistency is crucial for creating a standardized experience of presbyopia, which could be particularly useful in educational or training contexts. Furthermore, the low variability in simulated near points across participants, including those with differing baseline vision (normal vision, myopia, and post-LASIK), highlights the simulation’s reliability in achieving the targeted effect.
The discrepancy between algorithmically determined and measured presbyopic near points can be attributed to measurement precision, distance calculation algorithms, and the effect of viewing angle on readings. Non-linear diopter adjustment and debounce algorithms may also contribute by moderating accommodation and filtering minor changes. Additionally, variations can arise from the interaction between the simulated effect and the participant’s natural accommodation ability.
While OpticalAging effectively simulates distance-dependent blur, it currently renders this blur uniformly. This contrasts with insights from our preliminary work, where some older adults described their close-range blur experience as involving “ghosting” or “double images” rather than simple diffusion. This suggests an area for future refinement to enhance simulation fidelity.
Furthermore, the fixed age modes (40s, 50s, 60s) based on population averages, while effective in demonstrating controlled near point shifts (Figure 7), may not fully capture the individual variability in presbyopia progression. This aligns with observations from our preliminary discussions with older adults, whose self-reported functional near points varied individually.
@@ 4.3. Part II: Subjective Experience of Simulated Presbyopia (50s Mode) During Daily Reading Tasks
To complement our quantitative measurements of shifted near points, we designed this qualitative assessment to evaluate the subjective experience of participants using our presbyopia simulation system in daily tasks. Participants are the same group as from Part I, they joined this part with a 5-minute break between. Regarding these task-based assessments, specifically the reading tasks, the simulation was set to the 50s mode. This decision was based on several key factors: (1) Progression of Presbyopia: By the 50s, presbyopia is typically well-established in most individuals (Truscott, 2009). (2) Significant Impact on Reading: The 50s typically mark a point where presbyopia significantly affects near vision tasks, especially reading. (3) Relevance to Target Users: Many users of presbyopia-simulating technology are likely to be younger individuals or designers seeking to understand the experience of more advanced presbyopia. (4) Comparative Baseline: For younger participants, this mode provides a clear contrast to their natural vision, allowing for meaningful comparisons and insights into the challenges faced by individuals with presbyopia.
@@
After finishing their tasks, participants answered the questions using the 7-point Likert scale questionnaire to evaluate their comprehension of presbyopia following the use of the optical simulation method. The questionnaire consists of seven structured questions (Figure 9) and two open-ended questions: (1) What aspect of the presbyopia simulation did you find the most surprising or insightful? (2) What will you suggest to further improve this presbyopia simulation approach? The questions are designed to gather additional feedback on their experiences with the simulation system, offering qualitative insights. The responses to the open-ended questions were documented in voice recordings with the participant’s consent.
Part II: Results
Figure 9. This figure presents boxplots showing the distribution of participant responses to a baseline question (B) and six statements (Q1-Q6) about their experience with the presbyopia simulation. The baseline question (B) assessed participants’ initial understanding of presbyopia before the simulation. Q1-Q6 evaluated various aspects of the simulation experience after completing tasks with the 50s simulation mode. Responses were recorded on a 7-point Likert scale. Each box represents the interquartile range (IQR) with the median indicated by a vertical line. Whiskers extend to the minimum and maximum values within 1.5 times the IQR. Green triangles represent mean values. The red vertical lines indicate the median for each question. Key findings include: (B) low initial understanding of presbyopia (median = 2), (Q1) clear perceived difference in near vision focus (median = 6), (Q2) increased confidence in understanding presbyopia (median = 6), (Q3) enhanced empathy for individuals with presbyopia (median= 6), (Q4) improved ability to imagine presbyopia’s impact on daily activities (median = 7), (Q5) high likelihood of recommending the simulation (median = 7), and (Q6) potential influence on future design decisions for products used by older adults (median = 5, n = 11 participants with design backgrounds). The plot demonstrates a substantial shift from low initial understanding to high agreement across all aspects of the simulation experience.
Figure 9 presents participants’ responses to the baseline question (B) about their original understanding of presbyopia and the six questions (Q1-Q6) assessing various aspects of the simulation experience. Participants initially reported a low understanding of presbyopia (B: $M=2.37$, $median=2$), with responses concentrated between “No knowledge” and the midpoint of the scale.
The simulation experience yielded consistently positive responses across all measures: (1) Noticing Differences (Q1): Participants strongly agreed that they noticed clear differences in their ability to focus on near objects ($M=5.84$, $median=6$). (2) Increased Confidence (Q2): After the simulation, participants reported significantly higher confidence in their understanding of presbyopia ($M=5.95$, $median=6$). (3) Empathy Development (Q3): The simulation was highly effective in fostering empathy ($M=6.37$, $median=6$), with responses clustered at the upper end of the scale. (4) Imagining Daily Impact (Q4): Participants found it easy to imagine how presbyopia might affect daily activities ($M=6.47$, $median=7$), showing the highest mean score among all questions. (5) Recommendation Likelihood (Q5): There was a strong inclination to recommend the simulation experience to others ($M=6.16$, $median=7$), although with slightly more variability in responses. (6) Influence on Design Decisions (Q6): For the subset of participants involved in design related domains (n = 11), the experience showed potential to influence future design decisions ($M=5.85$, $median=5$), although with a greater spread of responses compared to other questions.
[…]
Eyesight Q1 Q2 Q3 Q4 Q5 Q6
Normal 5.33 (1.53) 5.67 (1.53) 6.00 (0.00) 6.67 (0.58) 6.00 (1.00) 7.00 (0.00)
Myopia 5.92 (1.24) 6.08 (0.79) 6.42 (0.67) 6.42 (0.67) 5.92 (1.00) 5.30 (1.06)
Post-LASIK 6.00 (1.41) 6.00 (0.82) 6.50 (0.58) 6.50 (0.58) 6.25 (1.50) 7.00 (0.00)
Table 3. Mean scores and standard deviations (in parentheses) for questionnaire items Q1-Q6, categorized by participants’ eyesight conditions. Scores are on a 7-point Likert scale, where higher values indicate stronger agreement. The eyesight categories include Normal vision, Myopia, and Post-LASIK. This table illustrates how participants with different visual backgrounds perceived various aspects of the presbyopia simulation experience.
Qualitative Feedback: Many were astonished by the level of vision impairment, especially the challenge of reading text at distances they once considered “far away”. Participants reported physical discomfort, such as hand fatigue, from maintaining extended reading positions. They expressed that “I feel sympathy for older individuals, my hands become sore after holding items away for a few minutes”, “It’s especially challenging because you need to hold things far away and keep your hands steady, which makes them sore, and it’s very frustrating when you really want to read but cannot. And since I’m nearsighted, it’s a completely different experience”. Several participants remarked on the difference between their normal vision and the simulated experience, particularly those with myopia. The experience also triggered emotional reactions, including sympathy, frustration, and anxiety about future vision changes, “for me, it (the experience) is very insightful, especially concerning my future. If it will be like that, then I’m worried”. Notably, participants displayed heightened curiosity about presbyopia, inquiring about its progression with age and possible prevention methods, “How can I prevent, or at least delay, presbyopia? I feel anxious about the fact that everyone will be affected by presbyopia”.
4.4. Part II: Discussion
The findings of this part, comprising both numerical data and detailed qualitative feedback, demonstrate the efficacy of our presbyopia simulation experience, providing a number of important insights:
  1. (1)
    Significant Knowledge Gain and Experiential Learning: The substantial increase in participants’ confidence coupled with participants’ reported astonishment at the level of vision impairment, suggests that the simulation serves as a powerful educational tool, effectively bridging the gap between abstract knowledge and embodied, experiential understanding. The qualitative data, such as participants’ astonishment at reading difficulties, strongly indicates a visceral learning experience that goes beyond simply being told about presbyopia’s effects. This aligns with theories of experiential learning (Kolb and Kolb, 2009) and supports the use of simulation in medical and design education (Okuda et al., 2009).
  2. (2)
    Enhanced Empathy and Perspective-Taking: The consistently high scores for empathy development ($M=6.37$) and ability to imagine daily impacts ($M=6.47$), along with their strong correlation ($r_{s}=0.72$), are further reinforced by the qualitative feedback. Participants reported physical discomfort and frustration, leading to increased sympathy for individuals with presbyopia. Comments like “I feel sympathy for older individuals, my hands become sore after holding items away for a few minutes” and “Now I understand why they tend to put things away while reading” demonstrate the simulation’s effectiveness in fostering emotional understanding and perspective-taking by providing a realistic presbyopia experience. This finding contributes to the growing body of research on empathy in design thinking (Leonard and Rayport, 1997) and supports the use of simulations in developing empathy among designers and healthcare professionals (Batt-Rawden et al., 2013).
  3. (3)
    Design Implications and Critical Thinking: The influence on future design decisions among the 11 participants with design backgrounds ($M=5.85$) is complemented by participants’ reflections on design considerations. For instance, one participant noted, “I started to think about how we should decide the sizes of the letters, which sizes are readable for users and which are not”. This indicates the simulation may prompt critical reflection on accessibility needs beyond simply influencing specific design choices. The moderate correlation between noticing differences in near focus and increased understanding confidence ($r_{s}=0.51$) further supports the idea that direct experience enhances comprehension, potentially leading to more informed design decisions.
  4. (4)
    Personal Relevance and Future Concerns: The qualitative feedback revealed an emotional dimension not captured in the quantitative data. Participants expressed anxiety about their own future vision changes, with comments like “If it will be like that, then I’m worried,” indicating a personal connection to the simulated experience. This personal relevance could be a powerful motivator for adopting inclusive design practices and could be further explored in future studies. It could also be considered as a sign of increased awareness of presbyopia.
  5. (5)
    Curiosity and Preventive Health Interest: The simulation sparked curiosity about presbyopia, with participants asking about its progression and prevention methods. This unexpected outcome suggests that the simulation could have broader implications for health education and preventive care awareness, extending its potential impact beyond the design field.
  6. (6)
    Varied Perceptions Based on Visual Background: While no statistically significant differences were found between eyesight categories, the trends observed (e.g., participants with vision correction experience reporting higher scores for noticing differences) suggest that personal visual history may influence the simulation experience. This underscores the importance of considering diverse user groups in both the development and application of such simulation tools.
While participants reported increased empathy, it is crucial to interpret this finding cautiously. As scholars like Bennett & Rosner (Bennett and Rosner, 2019) and Hartman (Hartman, 2025) note, empathy generated through simulation can be complex and may not fully capture or might even obscure aspects of lived experience. The reported increase in empathy might reflect a heightened awareness or appreciation of challenges, which, while positive, should primarily serve to motivate deeper engagement with users.

### v2

@@ 3.1. Replication and Generalizability
While the specific hardware platform used in our study is subject to a Non-Disclosure Agreement (NDA), our methodology and algorithmic principles are generalizable and designed for replication. The core components—commercially available tunable lenses and a distance sensor—can be integrated following the logic detailed in this paper.
The critical element for replication is the implementation of an algorithm that, based on distance input, applies a negative diopter adjustment to reduce the user’s accommodative amplitude to a desired, age-specific level (see Table 2), rather than a positive one for correction. To further aid replication, we have provided a detailed calibration and simulation workflow in our flowcharts (Appendix Figure 11 and Figure 4). We will also make our implementation notes and detailed algorithmic logic (e.g., pseudo-code) available as supplemental material to ensure the research community can build upon our work.
@@ 4. User Study
To explore and validate our approach, our user study was structured into three distinct parts. Part I involved the evaluation of optically shifted near points for simulated presbyopic age groups corresponding to the 40s, 50s, and 60s. Part II required participants to perform 20-minute daily reading activities with our simulation eyewear in the experimental environment. Finally, Part III was an exploratory N=2 case study with a professional product designer and a design student. The goal of this final part was not to produce generalizable findings, but to generate initial, qualitative insights into the practical application of our simulation and to illustrate its potential use within a professional design workflow. Our research protocols were rigorously reviewed and have obtained approval from the Institutional Review Board (IRB) of our university.
@@
Calibration and Individual Adjustment. The user study began with a careful calibration procedure to tailor to each participant’s unique visual needs. Using tunable lenses’ ability to simulate concave or convex lenses, nearsighted or farsighted participants participated in the study without their regular corrective eyewear. The corrective values determined by the participants were then incorporated into our simulation algorithm as the offset.
First, the participant puts on the ViXion01 eyewear and adjusts the interpupillary distance to position the tunable lens modules precisely in front of their eyes; see Figure 2 for a detailed hardware structure. Following ViXion01’s official guidelines, and our calibration workflow, see Figure 11, we ensure that the embedded tunable lenses initially functioned as corrective lenses to counteract potential nearsightedness or farsightedness. If the wearer has normal vision, the tunable lenses function as 0D lenses so that they can see things clearly as usual, when the focal plane is far away from the preset near point range. Participants who reported having no noticeable visual acuity difference between the two eyes are instructed to manipulate the adjustable dial, which changes the lenses’ diopters until clarity is achieved for a standing hardcover book at a distance of one meter. Individual lens adjustment is also performed for participants reporting anisometropia (disparate visual acuity between their eyes), Figure 11 demonstrates the detailed calibration workflow. In this individual adjustment case, all calibration steps were remotely monitored and controlled using ViXion01’s official smartphone application, the graphical user interface is illustrated in Figure 3 B.
@@ 4.1. Part I: Evaluation of Optically Shifted Near Points
Participants: We recruited 19 individuals from diverse professional backgrounds for this study via word of mouth. They ranged from 18 (with consent from the individual and their guardian) to 35 years old, with an average age (MEAN) of 27.8 years and a standard deviation (SD) of 4.1. We specifically targeted individuals under 40 years old, as those in their 40s typically begin to experience presbyopia (Duane, 1922). All participants were provided with essential information about this user study through a consent form. To ensure participant safety, we adopted exclusion criteria consistent with prior aging simulation studies (Lavallière et al., 2017). Participants were screened for history of dizziness, balance disorders, or recent falls. Additionally, the maximum continuous wear time was limited to 20 minutes to prevent disorientation or motion sickness. They took part in the study after signing the consent form. As for the baseline eyesight criteria, because the ViXion01’s built-in tunable lenses can optically correct both nearsightedness and farsightedness, participants were allowed to participate without wearing corrective glasses or contact lenses. Detailed calibration regarding the fitting of their interpupillary distance and corrective focal power is described in the previous Approach Section. Each participant received a 1000 JPY Amazon gift card as compensation for their time.
Figure 6. This plot includes data for the baseline (no simulation, natural near point) and simulated conditions representing optically shifted near point of 40s, 50s, and 60s for 19 participants. Each box represents the interquartile range (IQR) with the median indicated by the central line. Whiskers extend to the minimum and maximum values within 1.5 times the IQR. Outliers, marked with ‘x’, are data points beyond the whiskers. The plot demonstrates a clear progression in near point distances as the simulated age increases. Statistical analysis confirmed significant differences between all simulation modes ($p<.001$ for all pairwise comparisons).
@@
4.2. Part I Findings: Validation of Optically Shifted Near Points
The quantitative evaluation confirms that OpticalAging effectively and significantly modulates the user’s near point of accommodation. As shown in Figure 6, we observed a distinct, non-linear progression in near point distances corresponding to the simulated age increments. A repeated measures ANOVA (Greenhouse-Geisser corrected) revealed a significant main effect of simulation mode on near point measurements ($F(1.09,19.64)=2365.2,p<.001$) with an exceptionally high effect size ($\eta_{p}^{2}=0.992$). This magnitude of effect validates the system’s capability to enforce a highly controlled physiological constraint, effectively overwriting the user’s natural accommodation.
Post-hoc analyses (Bonferroni correction) confirmed significant differences between all condition pairs ($p<.001$). Specifically, the median near point shifted from a baseline of 132 mm to 233 mm (40s), 378 mm (50s), and 782 mm (60s). These measured values align closely with the theoretical presbyopia models defined by (Duane, 1922; Presbyopia, 2016; Schwartz, 2013) that informed our algorithm. Crucially, the low inter-participant variability—persisting across diverse baseline visual conditions (normal, myopia, post-LASIK)—demonstrates the system’s robustness in delivering a standardized impairment experience. While minor deviations from algorithmic targets were observed, likely attributable to the system’s debounce logic and measurement precision, the simulation successfully replicates the functional deficit required for design evaluation.
@@ 4.3. Part II: Subjective Experience of Simulated Presbyopia (50s Mode) During Daily Reading Tasks
To complement our quantitative measurements of shifted near points, we designed this qualitative assessment to evaluate the subjective experience of participants using our presbyopia simulation system in daily tasks. Participants were in the same group as in Part I, they joined this part after a 5-minute break. Regarding these task-based assessments, specifically the reading tasks, the simulation was set to the 50s mode. This decision was based on several key factors: (1) Progression of Presbyopia: By the 50s, presbyopia is typically well-established in most individuals (Truscott, 2009). (2) Significant Impact on Reading: The 50s typically mark a point where presbyopia significantly affects near vision tasks, especially reading.
(3) Relevance to Target Users: Many users of presbyopia-simulating technology are likely to be younger individuals or designers seeking to understand the experience of more advanced presbyopia.
(4) Comparative Baseline: For younger participants, this mode provides a clear contrast to their natural vision, allowing for meaningful comparisons and insights into the challenges faced by individuals with presbyopia.
@@
Figure 8. This figure presents boxplots showing the distribution of participant responses to a baseline question (B) and six statements (Q1-Q6) about their experience with the presbyopia simulation. The baseline question (B) assessed participants’ initial understanding of presbyopia before the simulation. Q1-Q6 evaluated various aspects of the simulation experience after completing tasks with the 50s simulation mode. Responses were recorded on a 7-point Likert scale. Each box represents the interquartile range (IQR) with the median indicated by a vertical line. Whiskers extend to the minimum and maximum values within 1.5 times the IQR. Green triangles represent mean values. The red vertical lines indicate the median for each question. Key findings include: (B) low initial understanding of presbyopia (median = 2), (Q1) clear perceived difference in near vision focus (median = 6), (Q2) increased confidence in understanding presbyopia (median = 6), (Q3) enhanced perspective-taking regarding individuals with presbyopia (median= 6), (Q4) improved ability to imagine presbyopia’s impact on daily activities (median = 7), (Q5) high likelihood of recommending the simulation (median = 7), and (Q6) potential influence on future design decisions for products used by older adults (median = 5, n = 11 participants with design backgrounds). The plot demonstrates a substantial shift from low initial understanding to high agreement across all aspects of the simulation experience.
@@
After finishing their tasks, participants answered the questions using the 7-point Likert scale questionnaire to evaluate their comprehension of presbyopia following the use of the optical simulation method. The questionnaire consists of seven structured questions (Figure 8) and two open-ended questions: (1) What aspect of the presbyopia simulation did you find the most surprising or insightful? (2) What would you suggest to further improve this presbyopia simulation approach? The questions are designed to gather additional feedback on their experiences with the simulation system, offering qualitative insights. The responses to the open-ended questions were documented in voice recordings with the participant’s consent.
4.4. Part II Findings: Perceived Understanding and Perspective Taking
The simulation significantly bridged the gap between abstract knowledge and embodied appreciation of presbyopic challenges. Participants initially reported low baseline knowledge (median =2=2). Post-simulation, a Wilcoxon signed-rank test confirmed a statistically significant increase in their confidence regarding understanding presbyopia (Z=3.84,p<.001,r=0.62Z=-3.84,p<.001,r=0.62). This quantitative shift was strongly corroborated by high scores in self-reported perspective-taking (Q3, median =6=6) and the ability to imagine daily life impacts (Q4, median =7=7), with a strong positive correlation between these two factors ($r_{s}=0.72,p<.001$).
Qualitatively, the simulation induced a visceral realization of the physical labor associated with presbyopia—a nuance often lost in static guidelines. Participants did not just report “blur”; they reported the physical fatigue of compensation. As one participant noted regarding the need to extend their reach:
“It’s especially challenging because you need to hold things far away and keep your hands steady, which makes them sore… I feel sympathy for older individuals [because] my hands become sore after holding items away for a few minutes.”
This physical discomfort translated into emotional frustration and anxiety about their own future aging, suggesting that the tool effectively facilitates perspective-taking through physiological constraint. However, we interpret these “empathy” scores cautiously; while the simulation successfully highlights immediate perceptual barriers and physical compensatory costs, it functions as a sensitizing tool rather than a replacement for the lived experience of long-term adaptation (Bennett and Rosner, 2019).
4.5. Part III: An Exploratory Case Study Evaluating Impact on Design Practice
While the previous parts of our study provided quantitative validation and subjective ratings, this final part is an exploratory case study with two participants (N=2). We must emphasize that due to the small sample size, the goal here is not to draw generalizable conclusions. Rather, the purpose is to provide a rich, qualitative illustration of how the OpticalAging simulation could be integrated into a design process and to generate initial hypotheses about its potential impact that can inform future, larger-scale research.